# Efficient and Scalable Graph Generation through Iterative Local Expansion

### 14 Dec 2023 v1

@@ 4 Experiments
We assess our model’s capability to generate graphs with node counts beyond the training distribution through extrapolation (creating larger graphs) and interpolation (varying sizes within observed ranges). We use two datasets of planar and tree graphs, each with 128 training graphs. Training graph sizes are uniformly sampled from $[32,64]$ for extrapolation and from $[32,64]\cup[128,160]$ for interpolation. Our evaluation involves generating graphs with 48 to 144 nodes, producing 32 graphs per size for validation and 40 for testing. We report the validity and uniqueness rates of generated graphs.
@@
Figure 3: Extrapolation/interpolation to out-of-distribution graph sizes. The training graph size range is visually represented as the shaded area.
@@ D.1 Cost Function Ablation Study
Local Variation 2.1 95.0 3.8 100 5.5
Random 3.8 85.0 6.1 100 7.1
@@ Appendix F Graph Neural Networks
Despite the rich literature on graph neural networks, addressing all these requirements simultaneously presents a considerable challenge. Message-passing GNNs, although linear in complexity with respect to the number of edges, are shown to be no more expressive than the 1-Weisfeiler-Lehman test (Xu et al., 2019). Additionally, handling fully connected graphs effectively is a hurdle for them that might impede the processing of locally dense expanded graphs (Definition 1). In contrast, higher-order GNNs, while being more expressive, come with elevated computational costs. Our investigation featured various architectures, including message-passing GNNs, a transformer-based model advocated by Vignac et al. (2023a), and PPGN (Maron et al., 2020). Our experimentation revealed that PPGN performed exceptionally well, although being computationally expensive.
@@ F.1 Provably Powerful Graph Networks
Maron et al. (2020) introduce PPGN, a permutation-equivariante graph neural network architecture. The PPGN possesses provable 3-WL expressivity, which is categorically stronger than message-passing models and operates on dense graph representations.
@@ F.2 Local PPGN
In devising our model, we recognized that the graphs of our interest are sparse, a common quality among many real-world graphs. Additionally, our expanded graphs (Definition 1) are locally dense with fully interconnected clusters but globally sparse. Hence, our objective is to construct a model that is expressively local but globally sparse. Our proposed Local PPGN which is comparable to the PPGN architecture for a fully connected graph and a message-passing GNN for a graph devoid of triangles, appears to strike a favorable balance between expressivity and complexity.
The fundamental operation of our Local PPGN layer is an edge-wise message-passing mechanism. It assumes an underlying directed graph $\overrightarrow{G}=({\mathcal{V}},\overrightarrow{{\mathcal{E}}})$, where ${\mathcal{V}}$ is the set of nodes and $\overrightarrow{{\mathcal{E}}}$ is the set of directed edges. The emebeddings ${\bm{h}}^{(i,j)}\in\mathbb{R}^{h}$ associated with the edges $(i,j)\in\overrightarrow{{\mathcal{E}}}$ are updated as follows:
@@
With $\bigoplus$ instantiated as a sum, $\phi$ and $\gamma$ with multi-layer perceptrons as $\phi({\bm{x}},{\bm{y}})=\mathrm{MLP}_{1}({\bm{x}})\odot\mathrm{MLP}_{2}({\bm{y}})$ and $\gamma({\bm{x}},{\bm{y}})=\mathrm{MLP}_{3}\left({\bm{x}}\mathbin{\|}{\bm{y}}\right)$, we recover the PPGN layer as delineated in Section F.1 for a fully connected graph. To discern this, one may compare the scheme 20 with our message-passing-like formulation of the PPGN layer in Equation 18.
When applied to a general graph, one interpretation of our layer’s effects is the application of the PPGN layer on each fully connected subgraph within the main graph. This was the initial impetus behind our design choice; we aimed to fuse the expressivity of PPGN with the efficiency of message-passing GNNs. Given that our graph generation method only introduces local modifications in each step, we theorize that this level of local expressivity is sufficient.
@@ F.3 Architectural Details
The aforementioned GIN comprises a sequence of message-passing layers, where each layer applies a multi-layer perceptron to the summed node embeddings of all adjacent nodes. Analogously to the Local PPGN, all intermediate embeddings from these layers are concatenated, subjected to a dropout operation with a probability of 0.1, and projected to $d_{\mathrm{SignNet}}$ dimensions using a linear layer.
@@ F.4 Graph Neural Network Architecture Ablation Study
Local PPGN 2.1 95.0 3.8 100 5.5
GINE 3.9 57.5 11.9 97.5 6.4
@@ Appendix H Complexity Analysis
Two observations are made to bound the sizes of the iterates $G_{l}$ in the sequence. Since expansion can only increase the number of nodes in the graph, no $G_{l}$ has more than $n$ nodes. For the number of edges, a similar statement cannot be made as a refinement step can remove arbitrarily many edges. Theoretically, a graph $G_{l}$ with $l>0$ could encompass more than $m$ edges. Nonetheless, this scenario is improbable for a trained model - provided sufficient training data and model capacity - as graph coarsening, according to Definition 3, can only decrease the number of edges. Consequently, the coarse graphs used for model training do not contain more edges than the dataset graphs. For the purpose of this analysis, we assume that the number of edges in all $G_{l}$ is asymptotically bounded by $m$.
@@
We next bound the complexity of generating an iterate $G_{l}$ in the sequence. For $l=L$, this involves instantiating a singleton graph and predicting the expansion vector ${\bm{v}}_{L}$, which can be done in constant time and using constant space. For all other instances where $0\leq l<L$, given the graph $G_{l+1}$ and the expansion vector ${\bm{v}}_{l+1}$, the graph $G_{l}$ and expansion vector ${\bm{v}}_{l}$ are derived by constructing the expansion $\tilde{G}_{l}=\tilde{G}(G_{l+1},{\bm{v}}_{l+1})$. This is followed by sampling ${\bm{v}}_{l}$ and ${\bm{e}}_{l}$ and subsequently constructing $G_{l}=G(\tilde{G}_{l},{\bm{e}}_{l})$. Let $v_{\text{max}}$ denote the maximum cluster size, which is $2$ for the edge contraction set family. This allows us to establish an upper bound on the number of edges in the expansion $\tilde{G}_{l}$. The edge set of $\tilde{G}_{l}$ comprises, at most, $n_{l+1}\frac{v_{\text{max}}(v_{\text{max}}-1)}{2}$ intracluster edges and a maximum of $m_{l+1}v_{\text{max}}^{2}$ cluster interconnecting edges. It is reasonable to assume that $v_{\text{max}}$ is constant, as the distribution of expansion sizes can only encompass a constant number of categories. Consequently, no expanded graph will asymptotically exceed $m$ edges. Sampling ${\bm{v}}_{l}$ and ${\bm{e}}_{l}$ is done by querying the denoising model a constant number of times. The complexity for this depends on the underlying graph neural network architecture. For message-passing models, this is linear in the number of nodes and edges in the graph. Our Local PPGN model also achieves this efficiency if the graph contains at most $\mathcal{O}(m)$ many triangles. Finally, we bound the complexity of obtaining the node embeddings for $G_{l}$. The first step involves calculating the $k$ main eigenvalues and eigenvectors of the graph Laplacian of $G_{l+1}$. By using the method suggested in Vishnoi (2013), this can be achieved with a complexity of $\tilde{\mathcal{O}}(km_{l+1})$, where $\tilde{\mathcal{O}}$ hides polylogarithmic factors. Computing the node embeddings from this using SignNet is done in $\mathcal{O}(km_{l+1})$ time and space. The replication of the embeddings for the expansion $\tilde{G}_{l}$ is linear relative to the number of nodes in $\tilde{G}_{l}$. Given that we select $k$ as a constant, the aggregate complexity for computing the node embeddings equates to $\tilde{\mathcal{O}}(n+m)$.
@@ I.2 Evaluation Metrics
We use the same evaluation metrics as Martinkus et al. (2022) to compare the performance of our model with other graph generative models. We report the maximum mean discrepancy (MMD) between the generated and the training graphs for the following graph properties: degree distribution, clustering coefficient, orbit counts, spectrum, and wavelet coefficients. As a reference, we compute these metrics for the training set and report the mean ratio across all of these metrics as a comprehensive indicator of statistical similarity. It is important to note that for the point cloud dataset, given its k-nearest neighbor structure, the degree MMD is identically zero; consequently, it is not incorporated into the mean ratio computation. Analogously, for the tree dataset, both the clustering coefficient and orbit counts are excluded from the ratio computation for the same reasons.
@@ I.3 Hyperparameters and Training setup
EDGE (Chen et al., 2023b) Number of diffusion steps 128 128 128 128 128
Batch size 16 8 32 12 2
@@
AAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 21 Feb 2024 v2

@@ 4 Experiments
We assess our model’s capability to generate graphs with node counts beyond the training distribution through extrapolation (creating larger graphs) and interpolation (varying sizes within observed ranges). We use a planar and a tree dataset, each comprising 128 training graphs with sizes uniformly sampled from $[32,64]$ for extrapolation and from $[32,64]\cup[128,160]$ for interpolation. Our evaluation involves generating graphs with 48 to 144 nodes, producing 32 graphs per size for validation and 40 for testing. We report the validity and uniqueness rates of generated graphs.
@@
Figure 3: Extrapolation and interpolation to out-of-distribution graph sizes. The shaded area represents the training size range.
@@ D.1 Cost Function Ablation Study
Local Variation 2.1 95.0 4.0 100 5.9
Random 3.9 85.0 6.2 100 8.2
@@ Appendix F Graph Neural Networks
Despite the rich literature on graph neural networks, addressing all these requirements simultaneously presents a considerable challenge. Message passing GNNs, although linear in complexity with respect to the number of edges, are shown to be no more expressive than the 1-Weisfeiler-Lehman test (Xu et al., 2019). Additionally, handling fully connected graphs effectively is a hurdle for them that might impede the processing of locally dense expanded graphs (Definition 1). In contrast, higher-order GNNs, while being more expressive, come with elevated computational costs. Our investigation featured various architectures, including message passing GNNs, a transformer-based model advocated by Vignac et al. (2023a), and PPGN (Maron et al., 2020). Our experimentation revealed that PPGN performed exceptionally well, although being computationally expensive.
@@ F.1 Provably Powerful Graph Networks
Maron et al. (2020) introduce PPGN, a permutation-equivariante graph neural network architecture. The PPGN possesses provable 3-WL expressivity, which is categorically stronger than message passing models and operates on dense graph representations.
@@ F.2 Local PPGN
In devising our model, we recognized that the graphs of our interest are sparse, a common quality among many real-world graphs. Additionally, our expanded graphs (Definition 1) are locally dense with fully interconnected clusters but globally sparse. Hence, our objective is to construct a model that is expressively local but globally sparse. Our proposed Local PPGN which is comparable to the PPGN architecture for a fully connected graph and a message passing GNN for a graph devoid of triangles, appears to strike a favorable balance between expressivity and complexity.
The fundamental operation of our Local PPGN layer is an edge-wise message passing mechanism. It assumes an underlying directed graph $\overrightarrow{G}=({\mathcal{V}},\overrightarrow{{\mathcal{E}}})$, where ${\mathcal{V}}$ is the set of nodes and $\overrightarrow{{\mathcal{E}}}$ is the set of directed edges. The emebeddings ${\bm{h}}^{(i,j)}\in\mathbb{R}^{h}$ associated with the edges $(i,j)\in\overrightarrow{{\mathcal{E}}}$ are updated as follows:
@@
With $\bigoplus$ instantiated as a sum, $\phi$ and $\gamma$ with multi-layer perceptrons as $\phi({\bm{x}},{\bm{y}})=\mathrm{MLP}_{1}({\bm{x}})\odot\mathrm{MLP}_{2}({\bm{y}})$ and $\gamma({\bm{x}},{\bm{y}})=\mathrm{MLP}_{3}\left({\bm{x}}\mathbin{\|}{\bm{y}}\right)$, we recover the PPGN layer as delineated in Section F.1 for a fully connected graph. To discern this, one may compare the scheme 20 with our message passing-like formulation of the PPGN layer in Equation 18.
When applied to a general graph, one interpretation of our layer’s effects is the application of the PPGN layer on each fully connected subgraph within the main graph. This was the initial impetus behind our design choice; we aimed to fuse the expressivity of PPGN with the efficiency of message passing GNNs. Given that our graph generation method only introduces local modifications in each step, we theorize that this level of local expressivity is sufficient.
@@ F.3 Architectural Details
The aforementioned GIN comprises a sequence of message passing layers, where each layer applies a multi-layer perceptron to the summed node embeddings of all adjacent nodes. Analogously to the Local PPGN, all intermediate embeddings from these layers are concatenated, subjected to a dropout operation with a probability of 0.1, and projected to $d_{\mathrm{SignNet}}$ dimensions using a linear layer.
@@ F.4 Graph Neural Network Architecture Ablation Study
Local PPGN 2.1 95.0 4.0 100 5.9
GINE 4.0 57.5 12.2 97.5 7.2
@@ Appendix H Complexity Analysis
Two observations are made to bound the sizes of the iterates $G_{l}$ in the sequence. Since expansion can only increase the number of nodes in the graph, no $G_{l}$ has more than $n$ nodes. For the number of edges, a similar statement cannot be made, as a refinement step can remove arbitrarily many edges. Theoretically, a graph $G_{l}$ with $l>0$ could encompass more than $m$ edges. Nevertheless, this scenario is improbable for a trained model - provided sufficient training data and model capacity - as graph coarsening, according to Definition 3, can only decrease the number of edges. Consequently, the coarse graphs used for model training do not contain more edges than the dataset graphs. For the purpose of this analysis, we assume that the number of edges in all $G_{l}$ is asymptotically bounded by $m$.
@@
Next, we bound the complexity of generating an iterate $G_{l}$ in the sequence. For $l=L$, this involves instantiating a singleton graph and predicting the expansion vector ${\bm{v}}_{L}$, which can be done in constant time and using constant space. For all other instances where $0\leq l<L$, given the graph $G_{l+1}$ and the expansion vector ${\bm{v}}_{l+1}$, the graph $G_{l}$ and expansion vector ${\bm{v}}_{l}$ are derived by constructing the expansion $\tilde{G}_{l}=\tilde{G}(G_{l+1},{\bm{v}}_{l+1})$. This is followed by sampling ${\bm{v}}_{l}$ and ${\bm{e}}_{l}$ and subsequently constructing $G_{l}=G(\tilde{G}_{l},{\bm{e}}_{l})$. Let $v_{\text{max}}$ denote the maximum cluster size, which is $2$ for the edge contraction set family. This allows us to establish an upper bound on the number of edges in the expansion $\tilde{G}_{l}$. The edge set of $\tilde{G}_{l}$ comprises, at most, $n_{l+1}\frac{v_{\text{max}}(v_{\text{max}}-1)}{2}$ intracluster edges and a maximum of $m_{l+1}v_{\text{max}}^{2}$ cluster interconnecting edges. It is reasonable to assume that $v_{\text{max}}$ is constant, as the distribution of expansion sizes can only encompass a constant number of categories. Consequently, no expanded graph will asymptotically exceed $m$ edges. Sampling ${\bm{v}}_{l}$ and ${\bm{e}}_{l}$ is done by querying the denoising model a constant number of times. The complexity for this depends on the underlying graph neural network architecture. For message passing models, this is linear in the number of nodes and edges in the graph. Our Local PPGN model also achieves this efficiency if the graph contains at most $\mathcal{O}(m)$ many triangles. Finally, we bound the complexity of obtaining the node embeddings for $G_{l}$. The first step involves calculating the $k$ main eigenvalues and eigenvectors of the graph Laplacian of $G_{l+1}$. By using the method suggested in Vishnoi (2013), this can be achieved with a complexity of $\tilde{\mathcal{O}}(km_{l+1})$, where $\tilde{\mathcal{O}}$ hides polylogarithmic factors. Computing the node embeddings from this using SignNet is done in $\mathcal{O}(km_{l+1})$ time and space. The replication of the embeddings for the expansion $\tilde{G}_{l}$ is linear relative to the number of nodes in $\tilde{G}_{l}$. Given that we select $k$ as a constant, the aggregate complexity for computing the node embeddings equates to $\tilde{\mathcal{O}}(n+m)$.
@@ I.2 Evaluation Metrics
We use the same evaluation metrics as Martinkus et al. (2022) to compare the performance of our model with other graph generative models. We report the maximum mean discrepancy (MMD) between the generated and the test graphs for the following graph properties: degree distribution, clustering coefficient, orbit counts, spectrum, and wavelet coefficients. As a reference, we compute these metrics for the training set and report the mean ratio across all of these metrics as a comprehensive indicator of statistical similarity. It is important to note that for the point cloud dataset, given its k-nearest neighbor structure, the degree MMD is identically zero; consequently, it is not incorporated into the mean ratio computation. Analogously, for the tree dataset, both the clustering coefficient and orbit counts are excluded from the ratio computation for the same reasons.
@@ I.3 Hyperparameters and Training setup
EDGE (Chen et al., 2023b) Number of diffusion steps 128 256 32 256 256
Batch size 16 8 16 8 2